OpenStudy (anonymous):

@-1/Sq Rt. of @ where @ = theta (sorry didnt know what else to put there for theta) So Im supposed to be using the power rule for differentiation. I am familiar with the rule I just don't know the algebra to get there.

zepdrix (zepdrix):

\[\frac{ \theta-1 }{ \sqrt{\theta} }\] So let's try to simplify this first :) We can rewrite this as 2 fractions. Divide each of the top terms by sqrt(@)

OpenStudy (anonymous):

hmmm like 2 seperate terms?

zepdrix (zepdrix):

mhm, like this :) \[\frac{ \theta }{ \sqrt \theta }-\frac{ 1 }{ \sqrt \theta }\]

zepdrix (zepdrix):

simplify before taking the derivative :)

OpenStudy (anonymous):

so take denominator to 1/2 power?

zepdrix (zepdrix):

sure, that's probably a good idea :D

OpenStudy (anonymous):

me and algebra haven't had the best relationship

zepdrix (zepdrix):

lolol that stinks c:  do these steps make sense?